# A Learnable Wavelet Transformer for Long-Short Equity Trading and Risk-Adjusted Return Optimization

### v2

@@ VII-A1 Data Split and Leakage Prevention
VII-A2 Implementation and Model Selection
@@
VII-B Model Performance
Among 14 candidate industries in Table III, we compute the industry-level annualized rate of return (ARR) of an intraday trading strategy using the training period only and retain industries with ARR $\geq 10\%$. The selected set is then fixed for all subsequent experiments to avoid leakage and to focus on sectors where intraday trading is potentially profitable and sufficiently liquid. This procedure selects six industries, and we focus on Renewable Energy and Retail Consumer Goods for detailed analysis, due to their having the lowest ARR.
TABLE III: Industry-level annualized rate of return (ARR) used for sector selection. Industries with ARR $\geq 10\%$ are retained for subsequent experiments.
@@ VII-B Model Performance
VII-C Hyperparameter Sensitivity Analysis
VII-C1 $\lambda_{\mathrm{roi}}$ in Overfitting Penalty
@@
To analyze its effect, we sweep
$\lambda_{\mathrm{roi}}\in\{0.0,0.1,0.3,0.5,0.7,1.0,1.2\},$
train the model under the same configuration for each setting, and record the final training ROI over the full 4-year horizon together with the corresponding val ROI and Sharpe ratio. Results are shown in Fig. 6.
@@
VII-C2 $\gamma$ in Low-guided High-frequency Injection
@@
Effect with Transformer backbones In contrast, a Transformer backbone forms a deep stack of residual blocks with Layer Normalization:
$X^{(\ell+1)}=\mathrm{LN}\bigl(X^{(\ell)}+F^{(\ell)}(X^{(\ell)})\bigr).$
A first-order approximation yields a layer Jacobian of the form
$J^{(\ell)}\approx I+\beta\,\tilde{J}^{(\ell)}$.
Thus, increasing $\beta$ raises the effective layer-wise gain, and the backward signal can grow through the product of Jacobians:
@@
Figure 7: Effect of the LGHI module $\beta$ on training loss and gradient norms for Transformer backbones. Curves correspond to $\beta=0.01$ (red), $\beta=0.1$ (black), $\beta=0.5$ (green), and $\beta=1.0$ (orange).
@@
VII-C3 $\lambda_{\mathrm{spec}}$ in Neural Wavelet Loss
@@
We sweep
$\lambda_{\mathrm{spec}}\in\{0,\ 0.3,\ 1,\ 3,\ 10,\ 30,\ 100\},$
keeping all other settings fixed, and evaluate validation ROI and Sharpe. Fig. 8 summarizes the results on Renewable Energy. We select $\lambda_{\mathrm{spec}}$ using validation performance to avoid test-set leakage, and report test results with the selected value in the main results section.
@@
VII-D Ablation Study
VII-D1 Effect of Learnable Wavelet Filters
We evaluate the proposed neural wavelet layer by replacing fixed filters in the same Transformer backbone. We compare three front-ends: (i) Neural Wavelet, which learns the filters jointly with the trading objective using Parseval, overlap, and pass-band regularizers; (ii) Classic Wavelet, a fixed wavelet transform; and (iii) Transformer only, which uses raw price features. As shown in Table IV, Neural Wavelet attains the best test ROI and Sharpe in almost all industries under matched capacity, and also yields smaller maximum draw-down and more stable performance across seeds, indicating that learnable, regularized wavelet filters extract more tradable signals than fixed transforms or no decomposition.
TABLE IV: Multi-seed, multi-industry comparison of Transformer-based backbones with different front-ends. Results are reported as mean $\pm$ std over ten seeds. Best results in each row are highlighted in red.
@@ VII-D1 Effect of Learnable Wavelet Filters
Biotechnology ROI 0.601 $\pm$ 0.034 0.323 $\pm$ 0.021 0.124 $\pm$ 0.066
Sharpe 1.695 $\pm$ 0.050 1.086 $\pm$ 0.048 0.511 $\pm$ 0.189
Semiconductors ROI 1.104 $\pm$ 0.053 0.416 $\pm$ 0.021 0.333 $\pm$ 0.021
Sharpe 2.555 $\pm$ 0.081 1.293 $\pm$ 0.071 1.134 $\pm$ 0.035
Renewable Energy ROI 0.423 $\pm$ 0.074 0.261 $\pm$ 0.033 0.169 $\pm$ 0.018
Sharpe 2.775 $\pm$ 0.365 1.723 $\pm$ 0.161 1.173 $\pm$ 0.113
Life Insurance ROI 0.185 $\pm$ 0.004 0.144 $\pm$ 0.004 0.120 $\pm$ 0.010
Sharpe 1.472 $\pm$ 0.039 1.206 $\pm$ 0.027 0.976 $\pm$ 0.086
Medical Devices ROI 0.669 $\pm$ 0.049 0.462 $\pm$ 0.067 0.289 $\pm$ 0.036
Sharpe 1.979 $\pm$ 0.167 1.475 $\pm$ 0.178 1.060 $\pm$ 0.125
Retail Consumer Goods ROI 0.659 $\pm$ 0.055 0.468 $\pm$ 0.143 0.317 $\pm$ 0.187
Sharpe 2.465 $\pm$ 0.293 1.850 $\pm$ 0.132 1.291 $\pm$ 0.185
Overall (avg.) ROI 0.607 $\pm$ 0.045 0.346 $\pm$ 0.048 0.225 $\pm$ 0.056
Sharpe 2.157 $\pm$ 0.166 1.439 $\pm$ 0.103 1.024 $\pm$ 0.122
VII-D2 Effect of Using Only a Single Frequency Branch
We evaluate whether both frequency branches are necessary by comparing the full model (Low+High) with two ablations: Low-Freq only and High-Freq only. Table V shows that removing either branch consistently hurts performance. Across most industries, Low-Freq only outperforms High-Freq only in ROI and Sharpe, indicating that slowly varying trends are more informative than high-frequency residuals alone. However, Low+High achieves the best average ROI/Sharpe with lower seed variance, suggesting that the high-frequency branch provides complementary short-term refinements, whereas relying solely on high-frequency signals reduces mean performance and increases training instability.
TABLE V: Effect of two frequency branches. Results are mean $\pm$ std over ten seeds. Best in each row is red.
@@ VII-D2 Effect of Using Only a Single Frequency Branch
Biotechnology ROI 0.601 $\pm$ 0.034 0.305 $\pm$ 0.028 0.133 $\pm$ 0.050
Sharpe 1.695 $\pm$ 0.050 0.902 $\pm$ 0.117 0.429 $\pm$ 0.185
Semiconductors ROI 1.104 $\pm$ 0.053 0.561 $\pm$ 0.052 0.244 $\pm$ 0.092
Sharpe 2.555 $\pm$ 0.081 1.360 $\pm$ 0.176 0.646 $\pm$ 0.279
Renewable Energy ROI 0.423 $\pm$ 0.074 0.225 $\pm$ 0.021 0.098 $\pm$ 0.037
Sharpe 2.775 $\pm$ 0.365 1.477 $\pm$ 0.191 0.702 $\pm$ 0.303
Life Insurance ROI 0.185 $\pm$ 0.004 0.094 $\pm$ 0.009 0.041 $\pm$ 0.015
Sharpe 1.472 $\pm$ 0.039 0.783 $\pm$ 0.101 0.372 $\pm$ 0.161
Medical Devices ROI 0.669 $\pm$ 0.049 0.340 $\pm$ 0.032 0.148 $\pm$ 0.056
Sharpe 1.979 $\pm$ 0.167 1.053 $\pm$ 0.136 0.501 $\pm$ 0.216
Retail Consumer Goods ROI 0.659 $\pm$ 0.055 0.334 $\pm$ 0.165 0.145 $\pm$ 0.286
Sharpe 2.465 $\pm$ 0.293 1.312 $\pm$ 0.161 0.624 $\pm$ 0.329
Overall (avg.) ROI 0.607 $\pm$ 0.045 0.310 $\pm$ 0.051 0.135 $\pm$ 0.089
Sharpe 2.157 $\pm$ 0.166 1.148 $\pm$ 0.147 0.546 $\pm$ 0.246
VII-D3 Low-guided High-frequency Injection
We further examine how the low and high frequency branches should be fused. In the proposed design, the low-frequency representation attends to the high-frequency branch via a LGHI block. As a comparison, we remove the LGHI module and simply concatenate the two branches followed by a linear projection, while keeping the overall model size comparable. The multi-industry, multi-seed results are reported in Table VI.
TABLE VI: LGHI module vs simple concatenation. Results are mean $\pm$ std over ten seeds. Best in each row is red.
@@ VII-D3 Low-guided High-frequency Injection
Biotechnology ROI 0.601 $\pm$ 0.034 0.203 $\pm$ 0.037
Sharpe 1.695 $\pm$ 0.050 0.640 $\pm$ 0.040
Semiconductors ROI 1.104 $\pm$ 0.053 0.374 $\pm$ 0.022
Sharpe 2.555 $\pm$ 0.081 0.964 $\pm$ 0.060
Renewable Energy ROI 0.423 $\pm$ 0.074 0.150 $\pm$ 0.035
Sharpe 2.775 $\pm$ 0.365 1.047 $\pm$ 0.165
Life Insurance ROI 0.185 $\pm$ 0.004 0.063 $\pm$ 0.012
Sharpe 1.472 $\pm$ 0.039 0.555 $\pm$ 0.034
Medical Devices ROI 0.669 $\pm$ 0.049 0.227 $\pm$ 0.028
Sharpe 1.979 $\pm$ 0.167 0.747 $\pm$ 0.076
Retail Consumer Goods ROI 0.659 $\pm$ 0.055 0.222 $\pm$ 0.154
Sharpe 2.465 $\pm$ 0.293 0.930 $\pm$ 0.130
Overall (avg.) ROI 0.607 $\pm$ 0.045 0.207 $\pm$ 0.048
Sharpe 2.157 $\pm$ 0.166 0.814 $\pm$ 0.083
@@
VII-D4 Sharpe Regularizer Ablation
To quantify the benefit of explicitly optimizing a trading-oriented objective, we ablate the Sharpe loss term by training two WaveLSFormer variants with identical architectures, optimization settings, and data splits: (i) Soft-label only and (ii) Soft-label with a Sharpe loss computed from per-step P&L. For each industry and seed, both variants share the same training schedule and differ only in whether the Sharpe term is included. Table VII reports industry-level test metrics averaged over 10 seeds with checkpoints selected by validation ROI. Adding the Sharpe loss improves risk-adjusted performance and often increases ROI while reducing maximum draw-down, suggesting that prediction-only training is insufficient and that directly regularizing the mean-volatility trade-off yields more favorable risk-return profiles.
@@ VII-D4 Sharpe Regularizer Ablation
$D_{t}=1-\frac{W_{t}}{\max_{s\leq t}W_{s}},\qquad\mathrm{MDD}=\max_{t}D_{t}.$
Figure 9: Equity curves with and without the Sharpe loss for Renewable Energy. Top: cumulative ROI on validation and test. Bottom: cumulative market return. The model without Sharpe loss (red) shows deeper and more frequent test draw-downs, whereas the Sharpe-aware model (orange) produces a smoother equity curve with smaller draw-downs at a similar final ROI, reflecting more conservative exposure in risky periods.
TABLE VII: Sharpe loss ablation by industry on the test set. Results are reported as mean $\pm$ standard deviation over 10 random seeds for ROI, Sharpe ratio and maximum draw-down (MDD, %).
@@
Overall (avg.) 0.470 $\pm$ 0.056 1.400 $\pm$ 0.224 15.463 $\pm$ 2.662 0.553 $\pm$ 0.035 2.000 $\pm$ 0.122 9.580 $\pm$ 1.736
VII-E Baseline Model Comparisons
VII-E1 Per-Industry Performance
@@
Table VIII reports test ROI and Sharpe by industry and on average. The wavelet front-end improves both metrics for every backbone, and LSTM-based models consistently outperform MLP, consistent with prior evidence that recurrent and attention architectures better capture long-range dependencies and regime shifts in financial time series [16, 17, 18, 24, 23]. Overall, WaveLSFormer achieves the best performance with a parameter budget comparable to LSTM variants, reaching $0.607\pm 0.045$ ROI and $2.157\pm 0.166$ Sharpe, compared with $0.225\pm 0.056$ ROI and $1.024\pm 0.122$ Sharpe for the plain Transformer. Fig. 10 shows consistent gains across industries, and Fig. 11 shows smoother equity growth with smaller draw-downs than the strongest non-wavelet baselines in Renewable Energy and Retail Consumer Goods.
TABLE VIII: Per-industry comparison of MLP, LSTM and Transformer backbones with and without the wavelet front-end. Results are mean $\pm$ std over ten seeds. The best in each row is red.
Industry Metric MLP Wav.+MLP LSTM Wav.+LSTM Trans. Wav.+Trans.
Biotechnology ROI $041\pm 0.013 0.0$$092\pm 0.024$ $0.106\pm 0.016$ $0.173\pm 0.026$ $0.124\pm 0.066$ $0.601\pm 0.034$
Sharpe $0.405\pm 0.157$ $0.553\pm 0.074$ $0.823\pm 0.242$ $0.934\pm 0.102$ $0.511\pm 0.189$ $1.695\pm 0.050$
Semiconductors ROI $0.110\pm 0.035$ $0.246\pm 0.064$ $0.285\pm 0.043$ $0.465\pm 0.069$ $0.333\pm 0.021$ $1.104\pm 0.053$
Sharpe $0.899\pm 0.349$ $1.227\pm 0.164$ $1.826\pm 0.537$ $2.072\pm 0.227$ $1.134\pm 0.035$ $2.555\pm 0.081$
Renewable Energy ROI $0.053\pm 0.020$ $0.139\pm 0.013$ $0.159\pm 0.017$ $0.221\pm 0.022$ $0.169\pm 0.018$ $0.423\pm 0.074$
Sharpe $0.851\pm 0.431$ $1.490\pm 0.054$ $1.482\pm 0.366$ $656\pm 0.061 1.1$$173\pm 0.113$ $2.775\pm 0.365$
Life Insurance ROI $040\pm 0.013 0.0$$089\pm 0.023$ $0.103\pm 0.015$ $0.168\pm 0.025$ $0.120\pm 0.010$ $0.185\pm 0.004$
Sharpe $0.774\pm 0.301$ $1.056\pm 0.141$ $1.572\pm 0.462$ $1.783\pm 0.195$ $0.976\pm 0.086$ $1.472\pm 0.039$
Medical Devices ROI $0.095\pm 0.030$ $0.214\pm 0.056$ $0.248\pm 0.037$ $0.404\pm 0.060$ $0.289\pm 0.036$ $0.669\pm 0.049$
Sharpe $0.840\pm 0.327$ $0.995\pm 0.133$ $1.707\pm 0.502$ $1.937\pm 0.212$ $1.060\pm 0.125$ $1.979\pm 0.167$
Retail Consumer Goods ROI $0.110\pm 0.028$ $0.208\pm 0.089$ $0.245\pm 0.047$ $0.471\pm 0.093$ $0.317\pm 0.187$ $0.659\pm 0.055$
Sharpe $110\pm 0.301 1.1$$154\pm 0.267$ $2.527\pm 0.862$ $2.895\pm 0.528$ $1.291\pm 0.185$ $2.465\pm 0.293$
Overall (avg.) ROI $0.075\pm 0.023$ $0.165\pm 0.045$ $0.191\pm 0.029$ $0.317\pm 0.049$ $0.225\pm 0.056$ $0.607\pm 0.045$
Sharpe $0.813\pm 0.311$ $1.079\pm 0.139$ $1.656\pm 0.495$ $1.879\pm 0.221$ $1.024\pm 0.122$ $2.157\pm 0.166$
@@
VII-E2 Model Complexity and Overall Performance
Table IX reports model complexity and trading performance for each backbone with and without the neural wavelet front-end. MLP is a lightweight reference with 8.1468.146M parameters and 3.4683.468G FLOPs, while LSTM and Transformer are substantially more expensive at 12.53812.538M/492.614492.614G and 15.92815.928M/665.921665.921G. To rule out brute-force scaling, we report parameters and FLOPs for paired comparisons within each backbone. The wavelet module introduces only small capacity changes, increasing parameters to 8.1518.151M for MLP, 13.29813.298M for LSTM, and 15.94315.943M for Transformer. Its computational overhead is modest, with LSTM FLOPs increasing by $7.8\%$ and Transformer FLOPs slightly decreasing by $0.9\%$.

### v3

@@ 7.1.1 Data Split and Leakage Prevention
7.1.2 Implementation and Model Selection
@@
7.2 Model Performance
Among 14 candidate industries in Table 3, we compute the industry-level annualized rate of return (ARR) of an intraday trading strategy using the training period only and retain industries with ARR $\geq 10\%$. The selected set is then fixed for all subsequent experiments to avoid leakage and to focus on sectors where intraday trading is potentially profitable and sufficiently liquid. This procedure selects six industries, and we focus on Renewable Energy and Retail Consumer Goods for detailed analysis, due to their having the lowest ARR.
Table 3: Industry-level annualized rate of return (ARR) used for sector selection. Industries with ARR $\geq 10\%$ are retained for subsequent experiments.
@@ 7.2 Model Performance
7.3 Hyperparameter Sensitivity Analysis
7.3.1 $\lambda_{\mathrm{roi}}$ in Overfitting Penalty
@@
To analyze its effect, we sweep $\lambda_{\mathrm{roi}}\in\{0.0,0.1,0.3,0.5,0.7,1.0,1.2\}$, and train the model under the same configuration for each setting, and record the final training ROI over the full 4-year horizon together with the corresponding val ROI and Sharpe ratio. Results are shown in Fig. 6.
@@
7.3.2 $\gamma$ in Low-guided High-frequency Injection
@@
Effect with Transformer backbones In contrast, a Transformer backbone forms a deep stack of residual blocks with Layer Normalization $X^{(\ell+1)}=\mathrm{LN}\bigl(X^{(\ell)}+F^{(\ell)}(X^{(\ell)})\bigr)$. A first-order approximation yields a layer Jacobian of the form $J^{(\ell)}\approx I+\beta\,\tilde{J}^{(\ell)}$. Thus, increasing $\beta$ raises the effective layer-wise gain, and the backward signal can grow through the product of Jacobians:
@@
Figure 7: Effect of the LGHI module $\beta$ on training loss and gradient norms for Transformer backbones.
@@
7.3.3 $\lambda_{\mathrm{spec}}$ in Neural Wavelet Loss
@@
We sweep $\lambda_{\mathrm{spec}}\in\{0,\ 0.3,\ 1,\ 3,\ 10,\ 30,\ 100\}$, keeping all other settings fixed, and evaluate validation ROI and Sharpe. Fig. 8 summarizes the results on Renewable Energy. We select $\lambda_{\mathrm{spec}}$ using validation performance to avoid test-set leakage, and report test results with the selected value in the main results section.
@@
7.4 Ablation Study
7.4.1 Effect of Learnable Wavelet Filters
We evaluate the proposed neural wavelet layer by replacing fixed filters in the same Transformer backbone. We compare three front-ends: (i) Neural Wavelet, which learns the filters jointly with the trading objective using Parseval, overlap, and pass-band regularizers; (ii) Classic Wavelet, a fixed wavelet transform; and (iii) Transformer only, which uses raw price features. As shown in Table 4, Neural Wavelet attains the best test ROI and Sharpe in almost all industries under matched capacity, and also yields smaller maximum draw-down and more stable performance across seeds, indicating that learnable, regularized wavelet filters extract more tradable signals than fixed transforms or no decomposition.
Table 4: Multi-seed, multi-industry comparison of Transformer-based backbones with different front-ends. Results are reported as mean $\pm$ std over ten seeds. Best results in each row are highlighted in bold.
@@ 7.4.1 Effect of Learnable Wavelet Filters
Biotechnology ROI $\bm{0.601\pm 0.034}$ 0.323 $\pm$ 0.021 0.124 $\pm$ 0.066
Sharpe $\bm{1.695\pm 0.050}$ 1.086 $\pm$ 0.048 0.511 $\pm$ 0.189
Semiconductors ROI $\bm{1.104\pm 0.053}$ 0.416 $\pm$ 0.021 0.333 $\pm$ 0.021
Sharpe $\bm{2.555\pm 0.081}$ 1.293 $\pm$ 0.071 1.134 $\pm$ 0.035
Renewable Energy ROI $\bm{0.423\pm 0.074}$ 0.261 $\pm$ 0.033 0.169 $\pm$ 0.018
Sharpe $\bm{2.775\pm 0.365}$ 1.723 $\pm$ 0.161 1.173 $\pm$ 0.113
Life Insurance ROI $\bm{0.185\pm 0.004}$ 0.144 $\pm$ 0.004 0.120 $\pm$ 0.010
Sharpe $\bm{1.472\pm 0.039}$ 1.206 $\pm$ 0.027 0.976 $\pm$ 0.086
Medical Devices ROI $\bm{0.669\pm 0.049}$ 0.462 $\pm$ 0.067 0.289 $\pm$ 0.036
Sharpe $\bm{1.979\pm 0.167}$ 1.475 $\pm$ 0.178 1.060 $\pm$ 0.125
Retail Consumer Goods ROI $\bm{0.659\pm 0.055}$ 0.468 $\pm$ 0.143 0.317 $\pm$ 0.187
Sharpe $\bm{2.465\pm 0.293}$ 1.850 $\pm$ 0.132 1.291 $\pm$ 0.185
Overall (avg.) ROI $\bm{0.607\pm 0.045}$ 0.346 $\pm$ 0.048 0.225 $\pm$ 0.056
Sharpe $\bm{2.157\pm 0.166}$ 1.439 $\pm$ 0.103 1.024 $\pm$ 0.122
7.4.2 Effect of Using Only a Single Frequency Branch
We evaluate whether both frequency branches are necessary by comparing the full model (Low+High) with two ablations: Low-Freq only and High-Freq only. Table 5 shows that removing either branch consistently hurts performance. Across most industries, Low-Freq only outperforms High-Freq only in ROI and Sharpe, indicating that slowly varying trends are more informative than high-frequency residuals alone. However, Low+High achieves the best average ROI/Sharpe with lower seed variance, suggesting that the high-frequency branch provides complementary short-term refinements, whereas relying solely on high-frequency signals reduces mean performance and increases training instability.
Table 5: Effect of two frequency branches. Results are mean $\pm$ std over ten seeds. Best in each row is bold.
@@ 7.4.2 Effect of Using Only a Single Frequency Branch
Biotechnology ROI $\bm{0.601\pm 0.034}$ 0.305 $\pm$ 0.028 0.133 $\pm$ 0.050
Sharpe $\bm{1.695\pm 0.050}$ 0.902 $\pm$ 0.117 0.429 $\pm$ 0.185
Semiconductors ROI $\bm{1.104\pm 0.053}$ 0.561 $\pm$ 0.052 0.244 $\pm$ 0.092
Sharpe $\bm{2.555\pm 0.081}$ 1.360 $\pm$ 0.176 0.646 $\pm$ 0.279
Renewable Energy ROI $\bm{0.423\pm 0.074}$ 0.225 $\pm$ 0.021 0.098 $\pm$ 0.037
Sharpe $\bm{2.775\pm 0.365}$ 1.477 $\pm$ 0.191 0.702 $\pm$ 0.303
Life Insurance ROI $\bm{0.185\pm 0.004}$ 0.094 $\pm$ 0.009 0.041 $\pm$ 0.015
Sharpe $\bm{1.472\pm 0.039}$ 0.783 $\pm$ 0.101 0.372 $\pm$ 0.161
Medical Devices ROI $\bm{0.669\pm 0.049}$ 0.340 $\pm$ 0.032 0.148 $\pm$ 0.056
Sharpe $\bm{1.979\pm 0.167}$ 1.053 $\pm$ 0.136 0.501 $\pm$ 0.216
Retail Consumer Goods ROI $\bm{0.659\pm 0.055}$ 0.334 $\pm$ 0.165 0.145 $\pm$ 0.286
Sharpe $\bm{2.465\pm 0.293}$ 1.312 $\pm$ 0.161 0.624 $\pm$ 0.329
Overall (avg.) ROI $\bm{0.607\pm 0.045}$ 0.310 $\pm$ 0.051 0.135 $\pm$ 0.089
Sharpe $\bm{2.157\pm 0.166}$ 1.148 $\pm$ 0.147 0.546 $\pm$ 0.246
7.4.3 Low-guided High-frequency Injection
We further examine how the low and high frequency branches should be fused. In the proposed design, the low-frequency representation attends to the high-frequency branch via a LGHI block. As a comparison, we remove the LGHI module and simply concatenate the two branches followed by a linear projection, while keeping the overall model size comparable. The multi-industry, multi-seed results are reported in Table 6.
Table 6: LGHI module vs simple concatenation. Results are mean $\pm$ std over ten seeds. Best in each row is bold.
@@ 7.4.3 Low-guided High-frequency Injection
Biotechnology ROI $\bm{0.601\pm 0.034}$ 0.203 $\pm$ 0.037
Sharpe $\bm{1.695\pm 0.050}$ 0.640 $\pm$ 0.040
Semiconductors ROI $\bm{1.104\pm 0.053}$ 0.374 $\pm$ 0.022
Sharpe $\bm{2.555\pm 0.081}$ 0.964 $\pm$ 0.060
Renewable Energy ROI $\bm{0.423\pm 0.074}$ 0.150 $\pm$ 0.035
Sharpe $\bm{2.775\pm 0.365}$ 1.047 $\pm$ 0.165
Life Insurance ROI $\bm{0.185\pm 0.004}$ 0.063 $\pm$ 0.012
Sharpe $\bm{1.472\pm 0.039}$ 0.555 $\pm$ 0.034
Medical Devices ROI $\bm{0.669\pm 0.049}$ 0.227 $\pm$ 0.028
Sharpe $\bm{1.979\pm 0.167}$ 0.747 $\pm$ 0.076
Retail Consumer Goods ROI $\bm{0.659\pm 0.055}$ 0.222 $\pm$ 0.154
Sharpe $\bm{2.465\pm 0.293}$ 0.930 $\pm$ 0.130
Overall (avg.) ROI $\bm{0.607\pm 0.045}$ 0.207 $\pm$ 0.048
Sharpe $\bm{2.157\pm 0.166}$ 0.814 $\pm$ 0.083
@@
7.4.4 Sharpe Regularizer Ablation
To quantify the benefit of explicitly optimizing a trading-oriented objective, we ablate the Sharpe loss term by training two WaveLSFormer variants with identical architectures, optimization settings, and data splits: (i) Soft-label only and (ii) Soft-label with a Sharpe loss computed from per-step P&L. For each industry and seed, both variants share the same training schedule and differ only in whether the Sharpe term is included. Table 7 reports industry-level test metrics averaged over 10 seeds with checkpoints selected by validation ROI. Adding the Sharpe loss improves risk-adjusted performance and often increases ROI while reducing maximum draw-down, suggesting that prediction-only training is insufficient and that directly regularizing the mean-volatility trade-off yields more favorable risk-return profiles.
@@ 7.4.4 Sharpe Regularizer Ablation
$D_{t}=1-\frac{W_{t}}{\max_{s\leq t}W_{s}},\quad\mathrm{MDD}=\max_{t}D_{t}.$
Figure 9: Equity curves with and without the Sharpe loss for Renewable Energy. Top: cumulative ROI on validation and test. Bottom: cumulative market return. The model without Sharpe loss shows deeper and more frequent test draw-downs, whereas the Sharpe-aware model produces a smoother equity curve with smaller draw-downs at a similar final ROI, reflecting more conservative exposure in risky periods.
Table 7: Sharpe loss ablation by industry on the test set. Results are reported as mean $\pm$ standard deviation over 10 random seeds for ROI, Sharpe ratio and maximum draw-down (MDD, %). The best overall results are in bold
@@
Overall (avg.) 0.470 $\pm$ 0.056 1.400 $\pm$ 0.224 15.463 $\pm$ 2.662 $\bm{0.553\pm 0.035}$ $\bm{2.000\pm 0.122}$ $\bm{9.580\pm 1.736}$
7.5 Baseline Model Comparisons
7.5.1 Per-Industry Performance
@@
Table 8 reports test ROI and Sharpe by industry and on average. The wavelet front-end improves both metrics for every backbone, and LSTM-based models consistently outperform MLP, consistent with prior evidence that recurrent and attention architectures better capture long-range dependencies and regime shifts in financial time series Nelson et al. (2017); Moghar and Hamiche (2020); Bao et al. (2017); Kabir et al. (2025); Mozaffari and Zhang (2024). Overall, WaveLSFormer achieves the best performance with a parameter budget comparable to LSTM variants, reaching $0.607\pm 0.045$ ROI and $2.157\pm 0.166$ Sharpe, compared with $0.225\pm 0.056$ ROI and $1.024\pm 0.122$ Sharpe for the plain Transformer. Fig. 10 shows consistent gains across industries, and Fig. 11 shows smoother equity growth with smaller draw-downs than the strongest non-wavelet baselines in Renewable Energy and Retail Consumer Goods.
@@
Table 8: Per-industry comparison of MLP, LSTM and Transformer backbones with and without the wavelet front-end. Results are mean $\pm$ std over ten seeds. The best results in each row are in bold.
Industry Metric MLP Wav.+MLP LSTM Wav.+LSTM Trans. Wav.+Trans.
Biotechnology ROI $041\pm 0.013 0.0$$092\pm 0.024$ $0.106\pm 0.016$ $0.173\pm 0.026$ $0.124\pm 0.066$ $\bm{0.601\pm 0.034}$
Sharpe $0.405\pm 0.157$ $0.553\pm 0.074$ $0.823\pm 0.242$ $0.934\pm 0.102$ $0.511\pm 0.189$ $\bm{1.695\pm 0.050}$
Semiconductors ROI $0.110\pm 0.035$ $0.246\pm 0.064$ $0.285\pm 0.043$ $0.465\pm 0.069$ $0.333\pm 0.021$ $\bm{1.104\pm 0.053}$
Sharpe $0.899\pm 0.349$ $1.227\pm 0.164$ $1.826\pm 0.537$ $2.072\pm 0.227$ $1.134\pm 0.035$ $\bm{2.555\pm 0.081}$
Renewable Energy ROI $0.053\pm 0.020$ $0.139\pm 0.013$ $0.159\pm 0.017$ $0.221\pm 0.022$ $0.169\pm 0.018$ $\bm{0.423\pm 0.074}$
Sharpe $0.851\pm 0.431$ $1.490\pm 0.054$ $1.482\pm 0.366$ $656\pm 0.061 1.1$$173\pm 0.113$ $\bm{2.775\pm 0.365}$
Life Insurance ROI $040\pm 0.013 0.0$$089\pm 0.023$ $0.103\pm 0.015$ $0.168\pm 0.025$ $0.120\pm 0.010$ $\bm{0.185\pm 0.004}$
Sharpe $0.774\pm 0.301$ $1.056\pm 0.141$ $1.572\pm 0.462$ $\bm{1.783\pm 0.195}$ $0.976\pm 0.086$ $1.472\pm 0.039$
Medical Devices ROI $0.095\pm 0.030$ $0.214\pm 0.056$ $0.248\pm 0.037$ $0.404\pm 0.060$ $0.289\pm 0.036$ $\bm{0.669\pm 0.049}$
Sharpe $0.840\pm 0.327$ $0.995\pm 0.133$ $1.707\pm 0.502$ $1.937\pm 0.212$ $1.060\pm 0.125$ $\bm{1.979\pm 0.167}$
Retail Consumer Goods ROI $0.110\pm 0.028$ $0.208\pm 0.089$ $0.245\pm 0.047$ $0.471\pm 0.093$ $0.317\pm 0.187$ $\bm{0.659\pm 0.055}$
Sharpe $110\pm 0.301 1.1$$154\pm 0.267$ $2.527\pm 0.862$ $\bm{2.895\pm 0.528}$ $1.291\pm 0.185$ $2.465\pm 0.293$
Overall (avg.) ROI $0.075\pm 0.023$ $0.165\pm 0.045$ $0.191\pm 0.029$ $0.317\pm 0.049$ $0.225\pm 0.056$ $\bm{0.607\pm 0.045}$
Sharpe $0.813\pm 0.311$ $1.079\pm 0.139$ $1.656\pm 0.495$ $1.879\pm 0.221$ $1.024\pm 0.122$ $\bm{2.157\pm 0.166}$
7.5.2 Model Complexity and Overall Performance
Table 9 reports model complexity and trading performance for each backbone with and without the neural wavelet front-end. MLP is a lightweight reference with 8.1468.146M parameters and 3.4683.468G FLOPs, while LSTM and Transformer are substantially more expensive at 12.53812.538M/492.614492.614G and 15.92815.928M/665.921665.921G. To rule out brute-force scaling, we report parameters and FLOPs for paired comparisons within each backbone. The wavelet module introduces only small capacity changes, increasing parameters to 8.1518.151M for MLP, 13.29813.298M for LSTM, and 15.94315.943M for Transformer. Its computational overhead is modest, with LSTM FLOPs increasing by $7.8\%$ and Transformer FLOPs slightly decreasing by $0.9\%$.